Illinois​ Furniture, Inc., produces all types of office furniture. The​ "Executive Secretary" is a chair that has been designed using ergonomics to provide comfort during long work hours. The chair sells for​ $130. There are

480480

minutes available during the​ day, and the average daily demand has been

5050

chairs. There are eight​ tasks:

                                                                                 

Task

Performance Time​(mins)

Task Must Follow Task Listed Below

A

44

minus−

B

77

minus−

C

77

​A, B

D

66

C

E

55

D

F

77

E

G

88

E

H

77

​F, G

e) For this​ process, the total idle time per cycle​ =

nothing

minutes ​(enter your response as a whole​ number).
Answer #1

Answer:

desired output = 50 units per 480 mins = 480secs /50= 9.6 mins per unit

Line balancing; the required cycle time is 9.6 mins per workstation. Sum of task times in a workstation cannot be more than cycle time.

Workstation

Tasks

Idle time= Cycle time- sum of task time

Reason

1

B

9.6-7= 2.6

We have to start with B, because B has longest processing time

2

A

9.6-4= 5.6

We cannot go foreword before completing A

3

C

9.6-7= 2.6

C is necessary before completing any other task

4

D

9.6-6= 3.6

5

E

9.6-5= 4.6

6

G

9.6-8= 1.6

Among F and G, G is the longest processing task

7

F

9.6-7= 2.6

8

H

9.6-7= 2.6

the total idle time = 2.6+ 5.6 +2.6+3.6+4.6+1.6+2.6+2.6= 25.8 mins

Answer e: Balanced delay = total idle time/ no. of workstation* cycle time = 25.8/(8*9.6)= 0.3359 ie 33.59%

Efficiency= 1-balanced delay= 66.40%